OpenStudy (anonymous):

Solve using the quadratic formula. x² - 9x + 20 = 0

OpenStudy (theviper):

Use\[\LARGE{-b \pm \sqrt{b^2-4ac} \over 2a}\]

OpenStudy (anonymous):

\[\begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} & {{\rm{when}}} & {ax^2 + bx + c = 0} \\ \end{array}\]

OpenStudy (anonymous):

can you tell us what a b and c are?

OpenStudy (theviper):

Here a=1, b=-9, c=20.

OpenStudy (klimenkov):

\[a=1,b=-9,c=20\]Solution is\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]Just put your numbers in the formula and count.

OpenStudy (anonymous):

b=9 c=20

OpenStudy (theviper):

@jeletta1 sorry b=-9

OpenStudy (theviper):

& wht about a

OpenStudy (anonymous):

a=1

OpenStudy (theviper):

yes

OpenStudy (theviper):

Can u solve it now???